OpenStudy (anonymous):

A .20 kg ball is moving at a speed of 12m/s. what is the momentum of the mall in kg times m/s?

OpenStudy (anonymous):

well... momentum is mass times velocity so 0.2 kg * 12 m/s = 2.4 kg m/s

OpenStudy (anonymous):

we know that momentum(p)=mass*velocity =m*v given:-mass=0.00002g velocity=12m/sec momentum=.00002*12 =.000024gram*meter/sec

OpenStudy (anonymous):

Be more careful - 0.2 kg = 200 g, not 0.00002 g \[0.2 kg* 10^{3}g/kg = 200g\]